# FlatQuant: Flatness Matters for LLM Quantization

### v1

@@ 1 Introduction
In this work, we introduce a new post-training quantization approach named FlatQuant (Fast and Learnable Affine Transformation). Our approach is grounded in the principle of achieving a flatter distribution of weights and activations, which is crucial for quantization. FlatQuant aims to identify the optimal affine transformation for each linear layer, employing a lightweight, block-wise training strategy over the calibration data. To minimize the inference overhead associated with affine transformations, FlatQuant harnesses the efficiency of Kronecker decomposition, thus reducing both the memory and computational demands. The proposed approach is compatible with various quantization techniques such as learnable clipping, and can be applied to various quantization settings, e.g., weight-only quantization or KV cache quantization. Additionally, by observing that affine transformations in FlatQuant are memory bound, we further fuse the affine transformations and quantization into a single kernel, thereby minimizing the global memory access and kernel lunch overhead. Lastly, extensive experiments are conducted to assess FlatQuant across various tasks, including language modeling and question answering, using LLaMA-2/3 models ranging from 7B to 70B parameters. The empirical results demonstrate that our proposed approach surpasses current state-of-the-art methods in terms of both accuracy and inference latency.
@@ 2.2 The Flatness for Quantization
We examine existing pre-quantization transformations with a focus on their potential for flatness, a critical factor for effective quantization. Intuitively, by removing outliers, these transformations are expected to produce flat weights and activations that are conducive to quantization. Additionally, the quantization error propagated through the network is also expected to be low and flat. However, our empirical results indicate that current methods are limited in achieving the desired flatness. This is in contrast to our proposed FlatQuant, which will be introduced in Section 3.
@@ 3.1 Fast and Learnable Affine Transformation
We decompose the original $\mathbf{P}\in\mathbb{R}^{n\times n}$ into $\mathbf{P}=\mathbf{P}_{1}\otimes\mathbf{P}_{2}$, where $\mathbf{P}_{1}\in\mathbb{R}^{n_{1}\times n_{1}},\mathbf{P}_{2}\in\mathbb{R}^{n_{2}\times n_{2}}$ are invertible matrices in smaller sizes, and $n=n_{1}n_{2}$. Recall the vectorization trick of the Kronecker product, i.e., $\textrm{vec}(\mathbf{V})(\mathbf{P}_{1}\otimes\mathbf{P}_{2})=\textrm{vec}(\mathbf{P}_{1}^{\top}\mathbf{V}\mathbf{P}_{2})$ for some $\mathbf{V}\in\mathbb{R}^{n_{1}\times n_{2}}$, the matrix multiplication in Equation 2 can be re-written as
@@
where $\tilde{\mathbf{X}}\in\mathbb{R}^{k\times n_{1}\times n_{2}}$ and $\tilde{\mathbf{W}}\in\mathbb{R}^{m\times n_{1}\times n_{2}}$ are reshaped from $\mathbf{X}$ and $\mathbf{W}$ accordingly, and ×$\times_{i}$ denotes the reduction over the $i$-th axis. Note that both weights and activations are converted back to matrix before multiplication. Such decomposition can save the memory up to $n/2$ times, given that $\frac{n^{2}}{n_{1}^{2}+n_{2}^{2}}\leq\frac{n^{2}}{2n_{1}n_{2}}=\frac{n}{2}$, with the equality holds when $n_{1}=n_{2}=\sqrt{n}$. Moreover, the computation saving is $\sqrt{n}/2$ times with the same optimal condition. In practice, we select $n_{1}^{*},n_{2}^{*}=\arg\min(n_{1}+n_{2})$, s.t. $n_{1}n_{2}=n$ and $n_{1}\leq n_{2}$. For instance, the optimal configuration is $(n_{1}^{*},n_{2}^{*})=(64,128)$ for $n=8192$. We find such a strategy gives the best speedup without compromising performance, as will be detailed in Figure 6.
@@ 3.2 Integration with the Transformer Architecture
We illustrate the integration of FlatQuant with a Transformer block based on an LLaMA-like architecture, as depicted in Figure 3. Following the conventional practices, we employ low-bit matrix multiplications for all linear layers, while keeping layer normalization layers, pre-quantization transformations, RoPE embeddings, and attention scores in FP16.
@@ 4.2 Main Results
We evaluate FlatQuant on MT-Bench using the LLaMA-3.1-8B-Instruct model in Table 3. While FlatQuant trails behind the FP16 baseline in coding and STEM tasks, it consistently outperforms QuaRot with GPTQ across all categories, narrowing the gap between the quantized model and the FP16 baseline. Notably, for math problems, FlatQuant matches the FP16 baseline’s score, exceeding QuaRot by 1.9 points. More evaluations are provided in Appendix C.1.
@@ 4.3 Inference Latency
Figure 4 shows the prefill and decoding speedup of FlatQuant across different batch sizes, with 2048 and 256 tokens for prefill and decoding, respectively. With kernel fusion and INT4 tensor core, FlatQuant can achieve up to 2.30x speedup for prefill and 1.76x speedup for decoding under the batch size of 64. Notably, FlatQuant is apparently faster than QuaRot (Ashkboos et al., 2024) thanks to the Kronecker decomposition and efficient kernel design. Although there is still a minor gap compared to the vanilla INT4 quantization, it significantly enhances accuracy and facilitates the deployment of INT4 LLMs in real-world applications.
@@
We now investigate the impact of the five online transformations (i.e., $\{\mathbf{P}_{a},\mathbf{P}_{o},\mathbf{P}_{h},\mathbf{P}_{ug},\mathbf{P}_{d}\}$) in FlatQuant on the overall speedup, as shown in Figure 6. Even with five per-layer transformations, FlatQuant results in a minimal 0.07x end-to-end slowdown, significantly outperforming QuaRot’s 0.26x with just three Hadamard transformations. Specifically, FlatQuant’s $\mathbf{P}_{d}$ causes a 0.04x slowdown due to large FFN intermediate sizes, compared with QuaRot’s 0.17x. Meanwhile, $\mathbf{P}_{o}$ results in a 0.01x slowdown, versus QuaRot’s 0.1x. The rest transformations (i.e., $\mathbf{P}_{a}$ and $\mathbf{P}_{ug}$) have an insignificant impact of less than 0.01x.
@@ Appendix A Related Work
Quantization is a crucial technique for reducing memory footprint and accelerating inference by employing fewer bits for storage and computation, especially for the application of LLMs. Unlike previous models, LLMs are shown to exhibit outliers in activation and massive outliers in pivot tokens (Wei et al., 2022; Dettmers et al., 2022; Liu et al., 2024a; Sun et al., 2024), which can severely degrade quantization accuracy. To eliminate the negative impact of outliers, pre-quantization transformations have been widely adopted in weight-activation quantization (Xiao et al., 2023; Wei et al., 2023; Shao et al., 2023; Ma et al., 2024; Ashkboos et al., 2024; Liu et al., 2024b) as well as in fully quantized training (Xi et al., 2023). Additionally, several weight-only quantization methods (Lin et al., 2023; Chee et al., 2024; Tseng et al., 2024) also incorporate pre-quantization transformations.
@@
Considering that per-channel scaling corresponds to the diagonal elements of the affine transformation matrix, AffineQuant (Ma et al., 2024) proposes learning the equivalent affine transformation. However, their approach focuses on learning full-size diagonally dominant matrices and employs a gradual mask optimization method, which may hinder the full potential of affine transformation in reducing quantization loss. Moreover, due to the formidable overhead associated with full-sized matrix multiplication, AffineQuant can only apply affine transformation to a small fraction of linear layers. In contrast, we employ fast and learnable affine transformations without these limitations, leading to substantial accuracy improvements and practical speedup.
@@ Appendix C Additional Experiments
C.1 More Experimental Results
@@ C.2 Additional Discussions
C.3 Experiment Details of Figure 6
@@
C.4 Additional Analyses of Inference Latency
@@ D.1 More Visualizations of Weight and Activation Distributions
We visualize the distribution envelopes of both original and transformed weights and activations on the LLaMA models in Figure 10-14. It can be observed that neither per-channel scaling nor Hadamard transformation can fully smooth out outlier channels to produce flatness, still leaving outlier channels, especially on activations. On the other hand, the affine transformation learned by FlatQuant can effectively produce flatter distributions for both weights and activations which are easier to quantize.
@@ D.2 More Visualizations of Quantization Error Landscapes
We visualize the quantization error landscapes of LLaMA models in Figure 2 and Figure 15-18. With the affine transformation to smooth outliers, FlatQuant can effectively suppress the quantization errors at pivot tokens and ease the quantization error propagation, leading to a flatter quantization error landscape compared with per-channel scaling and Hadamard transformation.
@@ Appendix E Limitations
Figure 15: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-2-7B. Figure 15(a)-15(c) plot the MSE surface of each method, while Figure 15(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 16: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-2-13B. Figure 16(a)-16(c) plot the MSE surface of each method, while Figure 16(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 17: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-2-70B. Figure 17(a)-17(c) plot the MSE surface of each method, while Figure 17(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 18: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-3-70B. Figure 18(a)-18(c) plot the MSE surface of each method, while Figure 18(d) overlays these surfaces by dividing each MSE with that of FlatQuant.

### v2

@@ 1 Introduction
In this work, we introduce FlatQuant, a new post-training quantization approach. The name has dual significance: it stands for the proposed method (Fast and Learnable Affine Transformation) and emphasizes its goal of achieving flatter distributions of weights and activations, which are crucial for effective quantization. FlatQuant aims to identify the optimal affine transformation for each linear layer, employing a lightweight, block-wise training strategy over the calibration data. To minimize the inference overhead associated with affine transformations, FlatQuant harnesses the efficiency of Kronecker decomposition, reducing both the memory and computational demands. Our approach is compatible with various quantization techniques such as learnable clipping, and can be applied to various quantization settings, e.g., weight-only quantization or KV cache quantization. Additionally, as the affine transformations in FlatQuant are memory bound, we further fuse these transformations and quantization into a single kernel, thereby minimizing the global memory access and kernel lunch overhead. Lastly, extensive experiments are conducted to assess FlatQuant across various tasks, including language modeling and question answering, using LLaMA-2/3 models ranging from 7B to 70B parameters. The empirical results demonstrate that our proposed approach surpasses current state-of-the-art methods in terms of both accuracy and inference latency.
@@ 2.2 The Flatness for Quantization
We examine existing pre-quantization transformations with a focus on their potential for flatness, a critical factor for effective quantization. Intuitively, by removing outliers, these transformations are expected to produce flat weights and activations that are conducive to quantization. Additionally, the quantization error propagated through the network is also expected to be low and flat. However, our empirical results indicate that current methods are limited in achieving the desired flatness. This is in contrast to the proposed FlatQuant, as will be elaborated in Section 3. We leave how FlatQuant learns flatness in Appendix 8.
@@ 3.1 Fast and Learnable Affine Transformation
We propose to decompose the original $\mathbf{P}\in\mathbb{R}^{n\times n}$ into $\mathbf{P}=\mathbf{P}_{1}\otimes\mathbf{P}_{2}$, where $\mathbf{P}_{1}\in\mathbb{R}^{n_{1}\times n_{1}},\mathbf{P}_{2}\in\mathbb{R}^{n_{2}\times n_{2}}$ are invertible matrices in smaller sizes, and $n=n_{1}n_{2}$. Recall the vectorization trick of the Kronecker product, i.e., $\textrm{vec}(\mathbf{V})(\mathbf{P}_{1}\otimes\mathbf{P}_{2})=\textrm{vec}(\mathbf{P}_{1}^{\top}\mathbf{V}\mathbf{P}_{2})$ for some $\mathbf{V}\in\mathbb{R}^{n_{1}\times n_{2}}$, the matrix multiplication in Equation 2 can be re-written as
@@
where $\tilde{\mathbf{X}}\in\mathbb{R}^{k\times n_{1}\times n_{2}}$ and $\tilde{\mathbf{W}}\in\mathbb{R}^{m\times n_{1}\times n_{2}}$ are reshaped from $\mathbf{X}$ and $\mathbf{W}$ accordingly, and ×$\times_{i}$ denotes the reduction over the $i$-th axis. Note that both weights and activations are converted back to matrix before multiplication. Such decomposition can save the memory up to $n/2$ times, given that $\frac{n^{2}}{n_{1}^{2}+n_{2}^{2}}\leq\frac{n^{2}}{2n_{1}n_{2}}=\frac{n}{2}$, with the equality holds when $n_{1}=n_{2}=\sqrt{n}$. Moreover, the computation saving is $\sqrt{n}/2$ times with the same optimal condition. In practice, we select $n_{1}^{*},n_{2}^{*}=\arg\min(n_{1}+n_{2})$, s.t. $n_{1}n_{2}=n$ and $n_{1}\leq n_{2}$. For instance, the optimal configuration is $(n_{1}^{*},n_{2}^{*})=(64,128)$ for $n=8192$. We find such a strategy gives the best speedup without compromising performance, as will be detailed in Figure 6. The affine transformations are therein pretty light-weight: they only account for 2.61% of the overall computational FLOPs and 0.11MB extra memory when integrated with LLaMA-2-7B. More details of inference overhead are in Appendix C.5.
@@ 3.2 Integration with the Transformer Architecture
As depicted in Figure 3, we now illustrate the integration of FlatQuant with a Transformer block based on an LLaMA-like architecture. Following the conventional practices, we employ low-bit matrix multiplications for all linear layers, while keeping layer normalization layers, pre-quantization transformations, RoPE embeddings, and attention scores in FP16.
@@ 4.2 Main Results
We evaluate FlatQuant on MT-Bench with LLaMA-3.1-8B-Instruct model in Table 3. While FlatQuant trails behind the FP16 baseline in coding and STEM, it consistently outperforms QuaRot with GPTQ across all categories, narrowing the gap between the quantized model and the FP16 baseline. Notably, for math problems, FlatQuant matches the FP16 baseline’s score, exceeding QuaRot by 1.9 points. More evaluations are provided in Appendix C.2.
@@ 4.3 Inference Latency
Figure 4 shows the prefill and decoding speedup of FlatQuant across different batch sizes, with 2048 and 256 tokens for prefill and decoding, respectively. It can be found that even without kernel fusion, FlatQuant acheives comparable speed-up with QuaRot, thanks to kronecker decomposition of affine transformations. With kernel fusion, FlatQuant can achieve up to 2.30x speedup for prefill and 1.76x speedup for decoding under the batch size of 64, which is apparently faster than QuaRot (Ashkboos et al., 2024). Although there is still a minor gap compared to the vanilla INT4 quantization, it significantly enhances accuracy and facilitates the deployment of INT4 LLMs in real-world applications.
@@
We now investigate the impact of the five online transformations (i.e., $\{\mathbf{P}_{a},\mathbf{P}_{o},\mathbf{P}_{h},\mathbf{P}_{ug},\mathbf{P}_{d}\}$) in FlatQuant on the overall speedup, as shown in Figure 6. Even with five per-layer transformations, FlatQuant results in a minimal 0.07x end-to-end slowdown, significantly outperforming QuaRot’s 0.26x with just three Hadamard transformations. Specifically, FlatQuant’s $\mathbf{P}_{d}$ causes a 0.04x slowdown due to large FFN intermediate sizes, compared with QuaRot’s 0.17x. Meanwhile, $\mathbf{P}_{o}$ results in a 0.01x slowdown, versus QuaRot’s 0.1x. The rest transformations (i.e., $\mathbf{P}_{a}$ and $\mathbf{P}_{ug}$) have an insignificant impact of less than 0.01x. Finally, it can be found that even without kernel fusion, the additional transformations in FlatQuant is still on par with QuaRot, thanks to Kronecker decomposition.
@@ Appendix A Related Work
Quantization is a crucial technique for reducing memory footprint and accelerating inference by employing fewer bits for storage and computation, especially for the application of LLMs. Unlike previous models, LLMs are shown to exhibit outliers in activation and massive outliers in pivot tokens (Wei et al., 2022; Dettmers et al., 2022; Liu et al., 2024a; Sun et al., 2024), which can severely degrade quantization accuracy. To eliminate the negative impact of outliers, pre-quantization transformations have been widely adopted in weight-activation quantization (Xiao et al., 2023; Wei et al., 2023; Shao et al., 2023; Ma et al., 2024; Ashkboos et al., 2024; Liu et al., 2024b) as well as in fully quantized training (Xi et al., 2023). Additionally, several weight-only quantization methods (Lin et al., 2023; Chee et al., 2024; Tseng et al., 2024) incorporate pre-quantization transformations. Searched or learnable clipping thresholds for weights or activations (Lin et al., 2023; Duanmu et al., 2024; Ashkboos et al., 2024; Liu et al., 2024b; Shao et al., 2023) are also explored to eliminate outliers.
@@
Considering that per-channel scaling corresponds to the diagonal elements of the affine transformation matrix, AffineQuant (Ma et al., 2024) proposes learning the equivalent affine transformation. However, their approach focuses on learning full-size diagonally dominant matrices and employs a gradual mask optimization method, which may hinder the full potential of affine transformation in reducing quantization loss. Moreover, due to the formidable overhead associated with full-size matrix multiplication, AffineQuant can only apply affine transformation to a small fraction of linear layers. In contrast, we employ fast and learnable affine transformations without these limitations, leading to substantial accuracy improvements and practical speedup.
@@ Appendix C Additional Experiments
C.1 FlatQuant Leads to Flatness
Figure 8: Flatness and mean squared quantization error (MSE) of different Transformer blocks in LLaMA-3-8B during FlatQuant’s training process. Flatness is calculated as the total flatness of all weights and activations within a Transformer block.
Here we show that with the training objective of Equation 4, the learned affine transformations indeed flatten the distribution of weights and activations, as visualized in Figure 1 and Figure 2.
Quantifying the Flatness.
We quantitatively evaluate the flatness of weights and activations by analyzing their channel magnitude distributions. Specifically, each distribution is represented as a one-dimensional vector $\mathbf{d}$, as depicted in Figure 1. To measure flatness, we compute the mean squared error (MSE) between the observed distribution $\mathbf{d}$ and an idealized perfectly flat distribution $\mathbf{d}^{\prime}$. The flat distribution $\mathbf{d}^{\prime}$ is defined such that all channels possess equal magnitudes and the same $\ell_{2}$ norm as $\mathbf{d}$, i.e., $\mathbf{d}^{\prime}=\frac{\|\mathbf{d}\|_{2}}{\sqrt{N}}\cdot\mathbf{1}_{N}$, where $N$ is the number of channels and $\mathbf{1}_{N}$ is an $N$-dimensional vector with all entries equal to one. The MSE between $\mathbf{d}$ and $\mathbf{d}^{\prime}$ serves as a metric for flatness, with lower MSE values indicating a distribution that is closer to uniform across all channels.
With the above definition, we find that flatness well explains the optimization process of FlatQuant. In Figure 8, we plot the flatness and the training objective in Equation 4 (i.e., the MSE) of different Transformer blocks of LLaMA-3-8B during the training process. It can be observed that as the training loss decreases, the flatness also gets improved. This indicates that FlatQuantlearns better transformation to obtain a flatter distribution which eventually contributes to smaller quantization error, i.e. flatness matters for LLM quantization.
C.2 More Experimental Results
@@
Results on Qwen-2.5-Instruct.
To further validate the generality of FlatQuant, we conduct experiments on the Qwen-2.5-Instruct models. FlatQuantachieves near-lossless quantization (e.g., only 0.21% accuracy loss on QA tasks for Qwen-2.5-Instruct-32B). The results on language modeling and QA benchmarks are summarized in Table 11.
Model Method W Quantizer WikiText-2 C4 ARC-C ARC-E HellaSwag LAMBADA PIQA Winogrande Avg
7B FP16 - 8.36 14.37 51.37 75.80 79.57 67.61 80.20 69.93 70.75
FlatQuant RTN 8.46 13.94 51.71 77.69 78.42 57.46 76.93 69.53 68.62
32B FP16 - 5.32 10.45 58.62 77.02 85.25 75.14 81.39 73.16 75.10
QuaRot RTN 6.95 12.17 52.13 74.37 80.41 68.37 78.45 67.72 70.24
QuaRot GPTQ 6.54 11.65 56.06 76.52 81.83 71.26 78.78 69.06 72.25
FlatQuant RTN 5.80 10.86 58.62 78.58 83.72 75.26 80.74 72.45 74.89
Table 11: Evaluation results of FlatQuant on Qwen-2.5-Instruct models.
@@ C.3 Additional Discussions
C.4 Experiment Details of Figure 6
@@
C.5 Additional Analyses of Inference Latency
@@
Memory Consumption of Online Transformations.
We compute the parameter count of each online transformation below: (1) $\mathbf{P}_{a}$: $2(\sqrt{h_{d}})^{2}$; (2) $\mathbf{P}_{o}$: $a^{2}$; (3) $\mathbf{P}_{h}$: $(h_{d}/a)^{2}$; (4) $\mathbf{P}_{ug}$: $2(\sqrt{h_{d}})^{2}$; (5) $\mathbf{P}_{d}$: $2(\sqrt{h_{i}})^{2}$. The total parameter count in one Transformer block is $4h_{d}+2h_{i}+a^{2}+(h_{d}/a)^{2}$. The additional memory consumption during inference is $2(4h_{d}+2h_{i}+a^{2}+(h_{d}/a)^{2})$ bytes, which only consumes about 0.11MB extra memory space for LLaMA-2-7B.
@@ D.1 More Visualizations of Weight and Activation Distributions
We visualize the distribution envelopes of both original and transformed weights and activations on the LLaMA models in Figure 11-15. It can be observed that neither per-channel scaling nor Hadamard transformation can fully smooth out outlier channels to produce flatness, still leaving outlier channels, especially on activations. On the other hand, the affine transformation learned by FlatQuant can effectively produce flatter distributions for both weights and activations which are easier to quantize.
@@ D.2 More Visualizations of Quantization Error Landscapes
We visualize the quantization error landscapes of LLaMA models in Figure 2 and Figure 18-21. With the affine transformation to smooth outliers, FlatQuant can effectively suppress the quantization errors at pivot tokens and ease the quantization error propagation, leading to a flatter quantization error landscape compared with per-channel scaling and Hadamard transformation.
@@ Appendix E Limitations
Figure 18: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-2-7B. Figure 18(a)-18(c) plot the MSE surface of each method, while Figure 18(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 19: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-2-13B. Figure 19(a)-19(c) plot the MSE surface of each method, while Figure 19(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 20: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-2-70B. Figure 20(a)-20(c) plot the MSE surface of each method, while Figure 20(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 21: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-3-70B. Figure 21(a)-21(c) plot the MSE surface of each method, while Figure 21(d) overlays these surfaces by dividing each MSE with that of FlatQuant.